OpenStudy (sparklestaraa):

While in the United States visiting her grandmother, Kendra copied the famous family apple pie recipe. When she returns to England, she plans on making the same apple pie for her friends. To do this, Kendra needs to convert the flour from cups to grams. She knows that 0.5 cup of flour has a mass of 64 grams and 2 cups of flour has a mass of 256 grams.

OpenStudy (sparklestaraa):

A. Write a function that shows grams as a function of cups. Use g(x). B. Given this scenario, determine a reasonable domain for this function. C. Graph the relationship between the number of cups and grams of flour. D. Evaluate g(x) for any value of x within the domain. Interpret the meaning of your solution in the context of this problem. E. Look at the method you used to evaluate your function in part D. Explain another way you can evaluate the function for the same input. F. Interpret what it means to have g(-2), and determine if this is reasonable given the situation.

OpenStudy (sparklestaraa):

@undeadknight26 @DanJS

OpenStudy (sparklestaraa):

@butterflydreamer @texaschic101 @Nnesha

OpenStudy (sparklestaraa):

can someone please help?

OpenStudy (danjs):

yo

OpenStudy (danjs):

A. Write a function that shows grams as a function of cups. Use g(x).

OpenStudy (danjs):

She knows that 0.5 cup of flour has a mass of 64 grams and 2 cups of flour has a mass of 256 grams.

OpenStudy (danjs):

those are 2 points on a line

OpenStudy (sparklestaraa):

u want me to write the function?

OpenStudy (danjs):

(x,y) = (cups, grams)

OpenStudy (danjs):

(x,y) = (0.5 , 64) (x1,y1) = (2,256) slope?

OpenStudy (sparklestaraa):

ohh this is still A right?

OpenStudy (danjs):

m = (256 - 64) / (2 - 0.5) = 128

OpenStudy (danjs):

yeah you need to write the function, it is a line, and they gave you 2 points ( cups, grams)

OpenStudy (sparklestaraa):

so overall the function is what? with g(x)

OpenStudy (sparklestaraa):

is it g(x)= 128

OpenStudy (danjs):

y - y1 = m*(x-x1) use a point ( 2 , 256) y - 256 = 128*(x - 2) y - 256 = 128x - 256 y = g(x) = 128x The number of grams g(x) is 128 times the number of cups x g(x) = 128x

OpenStudy (danjs):

That's it. They gave you 2 points, ( x, g(x)) = (cups , grams ), just had to calculate the slope, and then make a line equation

OpenStudy (sparklestaraa):

:D OMG thanks :D it makes sense

OpenStudy (danjs):

could have recognized that it goes through the origin (0,0) , , 0 cups, is 0 grams. y = k*x (varies directly) k = 128 from using one of the points

OpenStudy (sparklestaraa):

this is B? @DanJS

OpenStudy (danjs):

That is the function g(x) = 128x for part A. Number of grams g(x) is 128 times #cups x

OpenStudy (danjs):

part b asks what are the reasonable values for x.

OpenStudy (danjs):

all the flour on earth, really, if you could get it all. lol

OpenStudy (sparklestaraa):

hhhh like all the tea in china xD

OpenStudy (danjs):

Reasonable domain would be , however many cups of flour you want to use , zero to whatever

OpenStudy (danjs):

just say like, the domain for g(x) is [0,20], x, the number of cups, is between 0 and 20 , it is arbitrary

OpenStudy (sparklestaraa):

oh ok thanks actually for u know simplifying things and saying just say xD it helps a lot :P plus its 10:57 so my mind is like what XD

OpenStudy (danjs):